INFRARED SPECTROSCOPY OF JET COOLED ND_{2}H_{2}^{+} MOLECULAR IONS: THE SYMMETRIC AND ANTISYMMETRIC NH STRETCH MODES

\underline{CHIH-HSUAN CHANG} and DAVID J NESBITT, JILA, National Institute of Standards and Technology University of Colorado, and Department of Chemistry and Biochemistry University of Colorado at Boulder, Colorado 80309.

Rovibrational progressions in the symmetric (v_{6}) and antisymmetric (v_{1}) NH stretching modes of the ND_{2}H_{2}^{+} molecular ion are observed for the first time, exploiting the i) high ion density and ii) high resolution capabilities of our slit jet discharge infrared spectrometer. These isotopomeric ions are generated by striking a modulated (50 KHz) electrical discharge in a mixture of ND_{3}/H_{2}O/H_{2} gases, achieving a modulated ion density suitable for time-gated, lock-in detection in the throat of a long path slit-jet expansion. Assignment of both b-type and c-type bands enables high accuracy determination of the rotational constants (A′′=4.85598(19), B′′=3.96811(11), and C′′=3.44661(40) cm^{−1}), with band origins for v_{1} and v_{6} modes determined to be 3297.54367(34) and 3337.90456(33) cm^{−1}, respectively. The results prove to be in good agreement with anharmonically corrected predictions from ab initio quartic force fields of Martin and Lee^{a}.

aJan. M. L. Martin and Timothy J. Lee, Chem. Phys. Lett. 258, 129 (1996)